A rotary vacuum filter is available with an area of 200 m2 and vacuum pressure of 75 kPa. Filter leaf tests have been performed on a cell broth with a viscosity of 5 cP. The leaf tests gave a specific cake resistance of 1×1011 cm/g and medium resistance of 1×108 cm-1. The cake solids (dry basis) per volume of filtrate was 15 g/L. It is desired to operate the large filter with cycle time of 45 s and a cake formation time of 10 s. What is the filtration rate expected for the rotary vacuum filter? How significant is the resistance medium?
1
Expert's answer
2020-07-30T07:10:39-0400

For rotatry filter we know that V=(2KAP)^1/2 where K=k/(1-)/C0 now A= area, P=pressure and k=cake permeability,